18 workers do a job in 40 days, if 6 workers were added from the beginning of the job , how many days it will need to finish that job?
Answer by stanbon(75887)

18 workers do a job in 40 days, if 6 workers were added from the beginning of the job , how many days will be needed to finish that job?
----
18*40 = 720 man-days are needed to do the job.
---
With 18+6 = 24 workers the time required would be 720/24 = 30
------------------
Another way to work it would be:
# of men and # of days are inversely related.
d = k/m
Solve for "k" using "18 men do the job in 40 days"
40 = k/18
k = 720
---
Equation:
d = 720/m
----
Then solve "if 6 workers were added from the beginning of the job , how many days will be needed to finish that job"?
---
d = 720/24
d = 30 days
==================
cheers,
Stan H.
============
--
Cheers,